OpenStudy (anonymous):

Simplify the expression. cos^2x+sin^2x/cot^2x-csc^2x

OpenStudy (ajprincess):

cot^2x-csc^2x=(cos^2x/sin^2x)-(1/sin^2x) =(cos^2x-1)/sin^2x =-(1-cos^2x)/sin^2x cos^2x+sin^2x=1 also 1-cos^2x=sin^2x. Does that help?

OpenStudy (anonymous):

yes kind of but if I simplify it even further what would the final answere be?

OpenStudy (anonymous):

would it be csc x? or -1 or 1 or sec x?

OpenStudy (ajprincess):

-(1-cos^2x)/sin^2x=?

OpenStudy (anonymous):

cot x?

OpenStudy (ajprincess):

nope 1-cos^2x=sin^2x so -(1-cos^2x)/sin^2x=-sin^2x/sin^2x=-1 cos^2x+sin^2x=1

OpenStudy (anonymous):

so its 1?

OpenStudy (ajprincess):

nope

OpenStudy (ajprincess):

check the sign

OpenStudy (anonymous):

-1?

OpenStudy (ajprincess):

yup

OpenStudy (anonymous):

ohhhh thank you so much:)

OpenStudy (ajprincess):

u r welcome:)

OpenStudy (anonymous):

can you help with this? Factor the algebraic expression below in terms of a single trigonometric function. csc 2x - 1

OpenStudy (ajprincess):

I am sry I'v gotta go nw. It's past midnight here

OpenStudy (anonymous):

ok gn:)

OpenStudy (ajprincess):

gn:)